# Lane Graph as Path: Continuity-preserving Path-wise Modeling for Online Lane Graph Construction

### v2

@@ 5.1 Dataset
We conduct experiments on two challenging and large-scale datasets, i.e., nuScenes [5] and Argoverse2 [41]. nuScenes [5] dataset consists of 1000 scene sequences. Each sequence is sampled in 2Hz frame rate and provides LiDAR point cloud and RGB images from 6 surrounding cameras, which covers ${}^{\circ}$ horizontal FOV of the ego-vehicle. The dataset provides the 2D lane graph without height information in the form of lane centerline and covers diverse online driving conditions (e.g., day, night, cloudy, rainy, and occlusion). For the online setting of lane graph construction, we set the perception ranges as $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis and preprocess the dataset following [20, 27, 23]. We train on the nuScenes train set and evaluate on the val set. The experiments on nuScenes dataset are conducted using 6 surrounding-view images by default. Argoverse2 [41] dataset contains 1000 scene logs and 7 surrounding cameras with ${}^{\circ}$ horizontal FOV. Each log is sampled in 10Hz frame rate and provides 3D lane graph. Due to the disability of pixel-wise modeling in performing 3D segmentation (piece-wise and path-wise modelings can simply predict an extra z corrdinate to support 3D lane graph construction), we drop the height information for fair comparison across different modelings. Other experimental settings are the same as nuScenes.
5.2 Baselines
@@
LaneGAP(ours) Path-wise 0.591 0.539 0.564 0.547 0.511 0.529 0.543 0.518 0.530 0.534 0.496 0.514 35.9M 16.5 15.6
@@
LaneGAP(ours) Path-wise 0.663 0.575 0.616 0.578 0.552 0.565 0.627 0.541 0.581 0.569 0.546 0.557 35.9M 13.8 13.0
@@
Implementation details. All the baselines use the same encoder consisting of ResNet50 [17] and GKT [10] to transform the ${}^{\circ}$ horizontal FOV multi-camera images to the unified BEV features at the resolution of $200\times 100$ (i.e., the BEV grid size is $0.3m$). All the models are controlled in comparable parameter sizes and trained for long enough epochs to ensure convergence (110 epochs on nuScenes and 24 epochs on Argoverse2). We trained all the experiments on 8 NVIDIA GeForce RTX 3090 GPUs with a total batch size of 32 (6 view images on nuScenes and 7 view images on Argoverse2). The only differences lie in the specialized decoders designed for respective modelings, which are detailed as follows.
@@
Pixel-wise modeling. We select HDMapNet [20] as the pixel-wise modeling baseline. It adopts a UNet structure with a ResNet18 to output the binarized segmentation map and direction map with two branches. Meanwhile, a cross-entropy loss is applied to the segmentation map, and $L_{2}$ loss is appled to the (-1,1) normalized direction map [18, 46]. To exclude the quantization error brought by classification on dense pixels, the output resolution of the UNet-shaped [36] segmentation model is set to $400\times 200$, which means the grid size is $0.15m$ (the perception range is $60m\times 30m$) the same as the graph interpolation size.
@@
Piece-wise modeling. STSU [6] is selected as the piece-wise modeling baseline, which consists of a detection branch to detect Bezier lane pieces with 4 control points and a association branch to predict the connectivity between lane pieces. It stacks 6 Transformer decoder layers to iteratively refine the predictions. $L_{1}$ loss is applied to the matched (0,1) normalized pieces and cross-entropy loss is applied to the connectivity. More recently, MapTR [23] proposes a much stronger Transformer decoder for detection branch. We further implement an additional piece-wise baseline by simply replacing the original Beizer piece detection branch of STSU with MapTR Polyline piece detection branch. We set the number of Polyline points to 20.
@@
Path-wise modeling. The path-wise LaneGAP adopts the same decoder as the piece-wise MapTR. We utilize deformable attention [47] as the cross attention of the Transformer and use the 30-point Polyline to represent the path and exploit the features along the path with resort to deformable attention as stated in Sec. 3.2. We stack 6 decoders and regress $(0,1)$ normalized coordinates. During training, $L_{1}$ loss is applied to the matched paths and cross-entropy loss is applied to the auxiliary BEV segmentation branch.
Table 2: Ablation for auxiliary BEV segmentation branch.
Table 3: Ablation for Polyline path modeling.
@@
Polyline Points Junction TOPO TOPO Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ 10 0.550 0.515 0.532 0.423 0.481 0.450 30 0.554 0.524 0.539 0.432 0.493 0.461 40 0.566 0.548 0.557 0.440 0.509 0.472
Bezier Points Junction TOPO TOPO Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ 3 0.480 0.483 0.481 0.309 0.447 0.365 5 0.514 0.507 0.510 0.375 0.477 0.477 10 0.498 0.503 0.501 0.334 0.463 0.388
Table 2: Ablation for auxiliary BEV segmentation branch.
Table 3: Ablation for Polyline path modeling.
Table 4: Ablation for Bezier path representation.
5.3 Quantitative Comparision
[…]
Highlights. Path-wise modeling achieves best $F_{1}$ score on the both subgraph around the junction points and the overall graph across two popular datasets (nuScenes and Argoverse2), while running at the fastest inference speed. The designed Path2Graph algorithm has negligible cost in translating predicted paths into directed lane graph (from 16.5 FPS to 15.6 FPS).
Path-wise vs. pixel-wise. Thanks to the high-resolution output of the segmentation model, pixel-wise modeling exhibits comparable precision for sub graph construction and overall graph construction. Nevertheless, the notorious fragmentation and oversmooth issues of segmentation methods make it hard to keep the continuity across pixels and distinguish the fine-grained subgraph around the junction points, leading to much lower recall compared to path-wise modeling. The sophiscated and prone-to-fail post-processing also adds enormous cost at inference to convert the segmentation output into vectorized lane graph. Compared to it, the path-wise modeling demonstrates higher lane graph construction quality and 2x faster inference speed on both nuScenes and Argoverse2 datasets.
Path-wise vs. piece-wise. Path-wise modeling outperforms piece-wise modeling on all metrics (precision, recall, parameter size, FPS). Even euipped with the same advanced Transformer decoder proposed in [23], our path-wise LaneGAP still significantly outperforms piece-wise MapTR, which validates the effectiveness of our proposed continuity-preserving modeling.
5.4 Qualitative Comparison
With the trained models in Tab. 1, Fig. 2 compares path-wise modeling with pixel-wise modeling and piece-wise modeling on complicated lane graphs with more than 4 junction points across different datasets.
@@ 5.4 Qualitative Comparison
Path-wise vs. pixel-wise. As shown in Fig. 2, the pixel-wise HDMapNet performs well on non-junction area. But for the sub-graph around the junction points, the segmentation model struggles to distinguish the fine-grained topology, and the post-processing is prone to fail to generate a decent vectorized lane graph, which aligns with the much lower Junction TOPO metrics in Tab. 1. Thanks to our rational modeling which treats each path as a whole for learning, we can capture the fine-grained topology with severe overlap and preserve the continuity of the lane.
Path-wise vs. piece-wise. The quality of lane graph construted by piece-wise modeling depends on the accuracy of both piece detection and connectivity prediction. As shown in Fig. 2, the piece-wise MapTR either detects jagged lane pieces or produces wrong connections, leading to incomplete and inaccurate lane graph after Piece2Graph post-processing. In contrast, our proposed path-wise modeling encodes the connectivity between pieces into the continuous path representation, which is easy to learn and robust.
@@ 5.5 Ablation Study
We ablate the design choices with a 24-epoch training schedule of path-wise modeling on nuScenes dataset by default without specification. And we report Junction TOPO and TOPO only on the directed lane graph.
@@
Polyline path representation. As shown in Tab. 4, the performance increases with adding more points for Polyline modeling under a 24-epoch schedule.
Bezier path representation. Tab. 4 shows that adding the number of control points from 3 to 5 improves 2.9% $F_{1}$ of Junction TOPO and 11.2% $F_{1}$ of TOPO, indicating too few Bezier control points cannot accurately describe the arbitrary shape in the broad perception. While adding the number of control points from 5 to 10 incurs an accuracy drop. Compared to Tab. 4, Bezier path representation is inferior to Polyline path representation, so we choose Polyline path representation as the default setting in LaneGAP.
@@
Modality Junction TOPO TOPO
@@
Vision-only 0.554 0.524 0.539 0.432 0.493 0.461
LiDAR-only 0.608 0.594 0.600 0.467 0.559 0.509
Vision & LiDAR 0.630 0.600 0.615 0.504 0.561 0.531
@@
6 Conclusion

### v3

@@ 5.1 Dataset
We conduct experiments on two popular and large-scale datasets, i.e., nuScenes [5] and Argoverse2 [47]. nuScenes [5] dataset consists of 1000 scene sequences. Each sequence is sampled in 2Hz frame rate and provides LiDAR point cloud and RGB images from 6 surrounding cameras, which covers 360 horizontal FOV of the ego-vehicle. The dataset provides the 2D lane graph without height information in the form of lane centerline and covers diverse online driving conditions. For the online setting of lane graph construction, we set the perception ranges as $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis, and preprocess the dataset following [22, 30, 26]. We train on the nuScenes train set and evaluate on the val set. The experiments on nuScenes dataset are conducted using 6 surrounding-view images by default. Argoverse2 [47] dataset contains 1000 scene logs and 7 surrounding cameras with 360 horizontal FOV. Each log is sampled in 10Hz frame rate and provides 3D lane graph. Due to the limitations of pixel-wise modeling in performing 3D segmentation (piece-wise and path-wise modelings can simply predict an extra z-coordinate to support 3D lane graph construction), we omit the height information for fair comparison across different modelings. Other experimental settings remain the same as those for nuScenes.
@@
LaneGAP Path-wise 0.591 0.539 0.564 0.547 0.511 0.529 0.543 0.518 0.530 0.534 0.496 0.514 35.9M 16.5 15.6
@@
LaneGAP Path-wise 0.663 0.575 0.616 0.578 0.552 0.565 0.627 0.541 0.581 0.569 0.546 0.557 35.9M 13.8 13.0
@@ 5.2 Baselines
Implementation details. All the baselines use the same encoder consisting of ResNet50 [18] and GKT [10] to transform the 360 horizontal FOV multi-camera images to BEV features at the resolution of $200\times 100$ (i.e., the BEV grid size is $0.3m$). All the models are controlled in comparable parameter sizes and trained for long enough epochs to ensure convergence (110 epochs on nuScenes and 24 epochs on Argoverse2). We train all the experiments on 8 NVIDIA GeForce RTX 3090 GPUs with a total batch size of 32 (6 view images on nuScenes and 7 view images on Argoverse2). Further details about training settings are provided in the supplementary material. The only differences lie in the specialized decoders designed for respective modelings, which are detailed as follows.
@@
Pixel-wise modeling. We select HDMapNet [22] as the pixel-wise modeling baseline. It adopts a U-Net structure with a ResNet18 to output a binarized segmentation map and a direction map with two branches. Meanwhile, a cross-entropy loss is applied to the segmentation map, and $L_{2}$ loss is applied to the (-1,1) normalized direction map [19, 54]. To exclude the quantization error brought by classification on dense pixels, the output resolution of the U-Net-shaped [41] segmentation model is set to $400\times 200$, which means the grid size is $0.15m$ (the perception range is $60m\times 30m$), the same as the graph interpolation size. The Pixel2Graph follows [22, 19].
@@
Piece-wise modeling. STSU [6] is selected as the piece-wise modeling baseline, which consists of a detection branch to detect Bezier lane pieces with 4 control points and an association branch to predict the connectivity between lane pieces. It stacks 6 Transformer decoder layers to iteratively refine the predictions. $L_{1}$ loss and Focal loss are applied to the matched (0,1) normalized pieces and cross-entropy loss is applied to the connectivity. More recently, MapTR [26] proposes a much stronger Transformer decoder for the detection branch. We further implement an additional piece-wise baseline by simply replacing the original Beizer piece detection branch of STSU with MapTR Polyline piece detection branch. We set the number of Polyline points to 20. The Piece2Graph follows [6].
@@
Path-wise modeling. The path-wise LaneGAP adopts the same decoder as the piece-wise MapTR. We use the 30-point Polyline to represent the path. We stack 6 deformable Transformer decoder layers and regress $(0,1)$ normalized coordinates. During training, $L_{1}$ loss and Focal loss are applied to the matched paths, and cross-entropy loss is applied to the auxiliary BEV segmentation branch. As for the Path2Graph, We discretize the path at the interval of $0.15m$, and the merging threshold is set to $0.15m$, the same as the graph interpolation size.
5.3 Quantitative Comparision
[…]
Highlights. Path-wise modeling achieves the best $F_{1}$ score on the both subgraph around the junction points and the overall graph across two popular datasets (nuScenes and Argoverse2), while running at the fastest inference speed. The designed Path2Graph algorithm has negligible cost in translating predicted paths into a directed lane graph (from 16.5 FPS to 15.6 FPS).
Path-wise vs. pixel-wise. Thanks to the high-resolution output of the segmentation model, pixel-wise modeling exhibits comparable precision for subgraph construction and overall graph construction. Nevertheless, the notorious fragmentation and oversmooth issues of segmentation methods make it hard to maintain continuity across pixels and distinguish the fine-grained subgraph around the junction points, leading to much lower recall compared to path-wise modeling. In comparison, path-wise modeling demonstrates higher lane graph construction quality and 2x faster inference speed.
Path-wise vs. piece-wise. Path-wise modeling outperforms piece-wise modeling on all metrics. Even equipped with the same advanced Transformer decoder proposed in [26], our path-wise LaneGAP still significantly outperforms piece-wise MapTR, which validates the effectiveness.
5.4 Quantitative Comparision on Openlane-V2 Dataset
[…]
5.5 Qualitative Comparison
Fig. 2 compares baselines on complicated lane graphs with more than 6 junction points across different datasets. Failure case studies and more qualitative comparisons are provided in the supplementary material.
@@ 5.5 Qualitative Comparison
Table 3: Auxiliary BEV segmentation branch.
Polyline Points Junction TOPO TOPO Param. FPS
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
10 0.550 0.515 0.532 0.423 0.481 0.450 35.9M 16.3
30 0.554 0.524 0.539 0.432 0.493 0.461 35.9M 15.6
40 0.566 0.548 0.557 0.440 0.509 0.472 35.9M 15.1
Table 4: Polyline path representation.
Bezier Points Junction TOPO TOPO Param. FPS
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
3 0.480 0.483 0.481 0.309 0.447 0.365 35.9M 16.1
5 0.514 0.507 0.510 0.375 0.477 0.477 35.9M 15.8
10 0.498 0.503 0.501 0.334 0.463 0.388 35.9M 15.6
Table 5: Bezier path representation.
Method L2 (m) $\downarrow$ Collision (%) $\downarrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
VAD-Tiny 0.46 0.76 1.12 0.78 0.21 0.35 0.58 0.38
VAD-Tiny+MapTR 0.43 0.74 1.06 0.74 0.20 0.31 0.51 0.34
VAD-Tiny+LaneGAP 0.42 0.70 0.99 0.70 0.18 0.27 0.43 0.29
Table 6: Comparison on downstream end-to-end planner VAD.
Method Modeling Junction TOPO TOPO
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
HDMapNet Pixel 0.401 0.328 0.361 0.415 0.349 0.379
STSU Piece 0.309 0.297 0.303 0.298 0.273 0.285
MapTR Piece 0.369 0.357 0.363 0.363 0.317 0.338
LaneGAP Path 0.384 0.400 0.392 0.384 0.392 0.388
Table 7: Comparing on large perception range $50m\times 100m$.
Method Modeling Junction TOPO TOPO
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
HDMapNet Pixel 0.464 0.312 0.373 0.468 0.318 0.379
STSU Piece 0.318 0.241 0.274 0.296 0.245 0.268
MapTR Piece 0.391 0.298 0.338 0.369 0.304 0.334
LaneGAP Path 0.458 0.356 0.400 0.447 0.358 0.398
Table 8: Comparing on new split.
HDMapNet STSU MapTR LaneGAP
mIoU on nusc 48.1 38.2 43.6 49.2
mIoU on av2 55.0 44.7 49.3 55.9
Table 9: Comparing on mIoU metric.
Post- Proc. Junction TOPO TOPO
Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
Pixel 0.898 0.678 0.773 0.913 0.693 0.788
Piece 1.000 1.000 1.000 1.000 1.000 1.000
Path 1.000 1.000 1.000 1.000 1.000 1.000
Table 10: Post-processing accuracy.
@@
Path-wise vs. pixel-wise. As shown in Fig. 2, the pixel-wise HDMapNet performs well on non-junction area. But for the subgraph around the junction points, the segmentation model struggles to distinguish the fine-grained topology, and the post-processing is prone to fail to generate a decent vectorized lane graph, aligning with the much lower Junction TOPO metrics in Tab. 1. On the contrary, LaneGAP can capture the fine-grained topology.
Path-wise vs. piece-wise. The quality of the lane graph constructed by piece-wise modeling depends on the accuracy of both piece detection and connectivity prediction. As shown in Fig. 2, the piece-wise MapTR either detects jagged lane pieces or produces wrong connections, leading to an incomplete and inaccurate lane graph after Piece2Graph post-processing. In contrast, our proposed path-wise modeling encodes the connectivity between pieces into the continuous path representation, which is easy to learn and robust.
@@ 5.6 Ablation Study
We ablate the design choices with a 24-epoch training schedule of path-wise modeling on nuScenes by default without specification. And we report Junction TOPO and TOPO only on the directed lane graph.
@@
Polyline path representation. As shown in Tab. 4, the accuracy increases with adding more points for Polyline modeling under a 24-epoch schedule, while the inference speed decreases. We choose 30-point Polyline as default setting.
Bezier path representation. Tab. 6 shows that adding the number of control points from 3 to 5 improves 2.9% $F_{1}$ of Junction TOPO and 11.2% $F_{1}$ of TOPO. While adding the number of control points from 5 to 10 incurs an accuracy drop.
Comparison on end-to-end planner VAD. Tab. 6 provides comparison on end-to-end planner VAD [21] by including an extra centerline lane graph branch. Since VAD utilizes instance-wise queries, HDMapNet is not applicable, we only compare with piece-wise MapTR. Path-wise modeling enhances planning performance and outperforms piece-wise modeling. Notably, LaneGAP’s path queries offer a complete depiction of the lane graph, unlike MapTR’s centerline piece queries, whose connectivity predictions are not directly usable in VAD.
Large perception range. In Tab. 8, we compare the performance on a large perception range ($50m\times 100m$) using the same training recipe of Tab. 1. LaneGAP still outperforms other methods.
Generalization on new split. StreamMapNet [52] proposes a new split of the nuScenes dataset, which splits the train/val set by the geolocation and reduces the map overlap. We further compare the performance on this new split in Tab. 8 using the same training recipe. LaneGAP maintains consistent superiority.
Comparison on mIoU metric. In Tab. 10, we further compare the performance of different modelings with mIoU metric following HDMapNet [22]. We render the prediction on the BEV map with $0.75m$ line width. The results further validate the superiority of LaneGAP.
Post-processing accuracy. In Tab. 10, we show the accuracy of the lane graphs transformed by respective post-processings given the ground-truth segmentation maps, sets of pieces and connections, and set of paths. The results indicate that Pixel2Graph post-processing is prone to fail in handling the fine-grained topology of the subgraph around junction points, as illustrated in Fig. 2.
6 Conclusion
Acknowledgement
This work was partly supported by the National Natural Science Foundation of China (NSFC 62276108).
Appendix 0.A Training Settings for Quantitative Comparison
In Tab. 12 and Tab. 12, we summarize the detailed training settings for the quantitative comparison in the main paper. We use the same settings under the long schedule.
Config Pixel-wise Piece-wise Path-wise
optimizer AdamW [35]
optimizer hyper-parameter $\beta_{1},\beta_{2},\epsilon=0.9,0.999,$ 1e-8
gradient clip norm type 2
gradient clip max norm 35
learning rate 6e-4
learning rate schedule cosine decay [34]
warmup iteration 500
weight decay 0.01
input resolution $6\times 900\times 1600$
image resize ratio 0.5
batch size 32
training epochs 110
BEV grid size $0.3m$
Table 11: Detailed training settings on nuScenes dataset. We use the same training settings for quantitative comparison.
Config Pixel-wise Piece-wise Path-wise
optimizer AdamW [35]
optimizer hyper-parameter $\beta_{1},\beta_{2},\epsilon=0.9,0.999,$ 1e-8
gradient clip norm type 2
gradient clip max norm 35
learning rate 6e-4
learning rate schedule cosine decay [34]
warmup iteration 500
weight decay 0.01
input resolution $7\times 2048\times 2048$
image resize ratio 0.3
batch size 32
training epochs 24
BEV grid size $0.3m$
Table 12: Detailed training settings on Argoverse2 dataset. We use the same training settings for quantitative comparison.
Appendix 0.B More Ablation Studies
@@
Modality Junction TOPO TOPO Param. FPS
@@
Vision-only 0.554 0.524 0.539 0.432 0.493 0.461 35.9M 15.6
LiDAR-only 0.608 0.594 0.600 0.467 0.559 0.509 9.1M 8.6
V & L 0.630 0.600 0.615 0.504 0.561 0.531 39.8M 5.5
@@
Tradeoff Ccomparison of modelings. In Tab. 15, we report the tradeoff comparison of modelings in terms of accuracy and cost on nuScenes dataset.
Method Modeling Modality $F_{1}$@Junction $F_{1}$@TOPO FPS Training time
HDMapNet Pixel-wise V 0.507 0.513 6.8 42h
MapTR Piece-wise V 0.496 0.465 15.0 49h
LaneGAP Path-wise V 0.564 0.529 15.6 58h
LaneGAP Path-wise V&L 0.632 0.597 5.5 60h
Table 15: Tradeoff comparison on nuScenes dataset. “V” and “L” respectively denote Vision modality and LiDAR modality. “$F_{1}$@Junction” and “$F_{1}$@TOPO” respectively denote $F_{1}$ of Junction TOPO and $F_{1}$ of TOPO on directed graph.
Appendix 0.C Visualizations for Ablations
In Fig. 4, we visualize the 24-epoch trained ablation experiments: multi-modality in column 2, LiDAR modality in column 3, vision modality in column 4 and 5, 30-point polyline representation in column 2, 3 and 4, 5-point Bezier representation in column 5. As shown in Fig. 4, the multi-modality method outputs a more accurate lane graph, and the Bezier method outputs smoother paths.
Figure 4: Visualizations for ablations. The visualized ablation experiments are under a 24-epoch training schedule. The multi-modality method provides more accurate lane graph, and the Bezier method outputs smoother paths.
Figure 5: Qualitative comparisons of increasing junction points. We qualitatively compare path-wise LaneGAP with pixel-wise HDMapNet and piece-wise MapTR on lane graphs with increasing junction points. All models perform well on simple lane graphs with few junction points. However, as the number of junction points increases, HDMapNet and MapTR struggle to predict reasonable lane graphs, while our proposed LaneGAP continues to deliver robust results.
Figure 6: Qualitative comparisons of diverse driving conditions. We qualitatively compare path-wise LaneGAP with pixel-wise HDMapNet and piece-wise MapTR across sunny, cloudy, rainy, dark, and occluded driving conditions. LaneGAP consistently demonstrates superior performance across all conditions. Specifically, in the occluded condition (row 5), results indicate that pixel-wise and piece-wise modelings, which may break continuity, are prone to predict incomplete lane graphs. In contrast, our path-wise modeling approach preserves the continuity of the lane graph, enhancing robustness in occluded driving conditions.
Figure 7: Qualitative comparisons of failure cases. We qualitatively compare path-wise LaneGAP with pixel-wise HDMapNet and piece-wise MapTR on the cases, where LaneGAP fails to handle. LaneGAP struggles to predict a complete subgraph for distant junction areas (row 1 and 2), located at the edge of the perception range. Furthermore, the extremely hard online driving conditions and topology (occlusion in row 4 and darkness in row 5) also make the construction very challenging. Addressing these issues requires further research and development.
Appendix 0.D More Qualitative Comparisons
Qualitative comparisons of increasing junction points. In Fig. 5, we compare path-wise LaneGAP with pixel-wise HDMapNet, and piece-wise MapTR on lane graphs with increasing junction points. All models perform well on simple lane graphs with few junction points. However, as the number of junction points increases, HDMapNet and MapTR struggle to predict reasonable lane graphs, while our proposed LaneGAP continues to deliver robust results.
Qualitative comparisons of diverse driving conditions. In Fig. 6, we compare these modelings across diverse driving conditions, including sunny, cloudy, rainy, dark, and occluded scenarios. LaneGAP consistently demonstrates superior performance across all conditions. Specifically, in the occluded condition (row 5), results indicate that pixel-wise and piece-wise modelings, which may break continuity, are prone to predict incomplete lane graphs. In contrast, our path-wise modeling approach preserves the continuity of the lane graph, enhancing robustness in occluded driving conditions.
Qualitative comparisons of failure cases. In Fig. 7, we compare these modelings on the cases, where LaneGAP fails to handle. The results show that LaneGAP struggles to predict a complete subgraph for distant junction areas (row 1 and 2), located at the edge of the perception range. Furthermore, the extremely hard online driving conditions and topology (occlusion in row 4 and darkness in row 5) also make the construction very challenging. Addressing these issues requires further research and development.
Appendix 0.E Video Visualizations
We visualize the whole nuScenes validation set and attach videos to the supplementary material. The videos are generated with two models: the path-wise method based on only vision input and the path-wise method based on vision and LiDAR input.